In a thin prism of glass of r.i 1.5 which of the following relations between the angle of minimum deviation(dm) and angle of refraction r will be correct? a)dm=r b)dm=1.5r c)dm=2r d)dm=r/2

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
dm = (\mu-1)* 2r = (1.5-1)*2r
dm = 0.5* 2r = r
Hence option A is correct.
 
Regards
Arun (askIITians forum expert)